how long does it take for $900 to double if it is invested at 7% compound continuously?
Answer by lwsshak3(11628)

how long does it take for $900 to double if it is invested at 7% compound continuously?
Formula for continuous compounding: A=Pe^rt, P=initial amt, r=interest rate, A=amt after t years
For given problem:
A/P=2
r=.07
e^rt=2
take log of both sides
rt*lne=ln2
lne=1
rt=ln2
t=ln2/r=ln2/.07≈9.9≈10 years
At 7% compound continuously,$900 to double in approximately 10 yrs